OpenStudy (anonymous):

solve for x and y: (1/4x-(1/2)=1 and (1/2)x-(1/3)y=12

OpenStudy (anonymous):

(1/4)x-(1/2)=1 (1/2)x-(1/3)y=12

OpenStudy (anonymous):

i hate fractions so i always get rid of them first x - 2 = 4 6x - 4y = 144 x = 6 6(6) - 4y = 144 36 - 4y = 144 -4y = 108 y = -27

OpenStudy (anonymous):

That can't be right, the answer sheet says it is (34,15)

OpenStudy (anonymous):

I plugged those into the equation and they worked. I can't figure out how to add the two equations to get that answer. As far as I got was 3x-2y=12 6x-4y=144 But I can't simplify them any further, because no matter what I multiply them by, the x and y both cancel out.

OpenStudy (anonymous):

your question is wrong then is it (1/4)x-(1/2)y=1?

OpenStudy (anonymous):

For the first one, yes. There is a second equation too.

OpenStudy (anonymous):

I forgot to write the y in my correction, my bad.

OpenStudy (anonymous):

x - 2y = 4 6x - 4y = 144 multiply first by -2 -2x + 4y = -8 6x - 4y = 144 add them 4x = 136 x = 34 x - 2y = 4 34 - 2y = 4 -2y = -30 y = 15 (34, 15)

OpenStudy (anonymous):

That's still not right...

OpenStudy (anonymous):

you just said the answer was 34,15 i'm completely lost as to what you are doing now.... how is my answer wrong?

OpenStudy (anonymous):

Wait, no, you are right. I just looked at it, and when I did my work on paper, I put (12/6)x as 3x instead of 2x, I see what happened. I'm sorry :D

OpenStudy (anonymous):

awesome ^_^